53.7   52.4   73.1   49.6   58.1   54.7   54.7   45.0   74.7   79.4   47.2   52.4   22.3   37.5   76.0   67.3   27.1   84.5   39.9   60.1   27.2   31.1   40.6   44.7   24.1   61.4   43.6   58.9   35.2   58.4 34.0   64.9   55.6   56.5   58.8   41.4   35.6   68.1   34.3   75.5

5. For each of the following values, use the one-tail five percent (5%) criteria and determine if the occurrence of the score would be considered relatively likely or unlikely? (Note: Assume that the researcher is only interested in the extreme low percents of students receiving free or reduced lunch--or the negative end of the normal curve). a. 67.3 b. 84.5 c. 34.0 6. For each of the following scores, use the two-tail five percent (5%) criteria and determine if the occurrence of the score would be considered relatively likely or unlikely? a. 43.6 b. 67.3
Answer #1

i)

mean = 51.49

sd = 16.2066

Z = (X - mean)/sd = (X - 51.49)/16.2066

a)

67.3

z= (67.3 - 51.49)/16.2066

= 0.9755284

critical value for one-tail five percent = 1.645

|z| < 1.645

hence likely

b)

84.5

z = (84.5 - 51.49)/16.2066

= 2.03682

|z | > 1.645

this is unlikely

c)

34

Z =(34 - 51.49)/16.2066

= - 1.07918

|Z| < 1.645

hence likely

ii)

a)

for two-tailed

critical values are 1.96 , -1.96

z = (43.6 - 51.49)/16.2066

= -0.48683

|Z| < 1.96

hence likely

b)

z = (67.3 - 51.49)/16.2066

= 0.97552

z < 1.96

relatively likely